A low-power laser used in a physics lab might have a power of 0.60 mW and a beam diameter of 3.3 mm.

(a) Calculate the average light intensity of the laser beam.
_________________ W/m2
(b) Compare it to the intensity of a light bulb producing 40 W light viewed from 2.4 m.
Ilaser / Ibulb = ___________
Answer #1

a) We have:

I = P/A = 0.60\times 10^{-3}/\pi(3.3\times 10 ^{-3}/2)^2 = 70.15 W/m^2

b) We have:

\frac{I_{laser}}{I_{bulb}} = \frac{70.15}{40/(4\pi(2.4)^2)} = 126.94